OpenStudy (anonymous):

how do you simplify a radical expression? like the square root of 27 or 150.

OpenStudy (anonymous):

√27 since 27 = 9 * 3: √(9 * 3) And now turn the square root of a product into the product of square roots: √9 * √3 Now I can solve for √9: 3 √3

OpenStudy (anonymous):

so if i simplified \[\sqrt{48}\] it would be \[2\sqrt{12}\] ?

OpenStudy (anonymous):

Almost. You take the radical, factor it and then take the sqrt of the factors: √48 = √(16×3) = √16√3 = 4√3

OpenStudy (mathstudent55):

In each case, you need to factor out the largest perfect-square factor there is. It is true that 48 = 4 * 12, but it is also 16 * 3.

OpenStudy (anonymous):

ohh i see. how would i simplify this expression? \[\sqrt{16/9}\]

OpenStudy (dape):

You have that \(16=4^2\) and \(9=3^2\), put that in and pull out the squares.

OpenStudy (anonymous):

Sqrt (16/9)= sqrt(16)/sqrt(9)= 4/3 ( I have a feeling I'm doing your homework for you :P )

OpenStudy (anonymous):

thanks. sorry I really don't know what i'm doing >.<

OpenStudy (anonymous):

Don't worry about it :D We're all here because we need some help. If you need anything else don't be afraid to ask around.